OpenStudy (anonymous):

please help me solve: 3a + 1 = -3.6(a - 1)

OpenStudy (anonymous):

All you have to do is bring the similar term aside of the '=' sign and solve it. Here, 3a +3.6a =3.6 -1 which becomes 6.6a =2.6 a = 2.6/6.6

OpenStudy (anonymous):

wait, so i add 3a + 3.6a together?

OpenStudy (saifoo.khan):

a = 13/33 simplified fraction.

OpenStudy (anonymous):

thanks! (: